An aluminum wire of diameter 4mm carries a current of 4.4 A.The current density is 

Grade:12

1 Answers

Arun
4 years ago

So here we've an aluminum wire of thickness 4mm

Thus, r=2mm and I=2.2A

Current Density,

J=I/A

A= πr²= π(2x10^-3

J= (2.2/π(2x10^-3)²)

 =1.75x0^5 A/m²